OpenStudy (anonymous):

int sin^5xcos^2x please help :)

OpenStudy (anonymous):

Trig sub, let u = ?

OpenStudy (anonymous):

duh nuh nuh duh nuh nuh batman!

OpenStudy (anonymous):

i know i have to separate the sin^5x

OpenStudy (anonymous):

so it would be int sin^4xsinxcos^2x

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

then would i replace the terms with the trig identities?

OpenStudy (caozeyuan):

reduction formula?

OpenStudy (anonymous):

i dunno :( lol

OpenStudy (xapproachesinfinity):

U sub^_^

OpenStudy (anonymous):

oh man it's one of the examples in the book :D :D

OpenStudy (anonymous):

this must be what heaven feels like... lol

OpenStudy (xapproachesinfinity):

first rewrite cos^5x using trig identities

OpenStudy (caozeyuan):

\[\int\limits \sin ^{5}x \cos ^{2}x dx=\int\limits \sin ^{4}x sinx cosx dx\]

OpenStudy (caozeyuan):

\[u=cosx\]\[du=-sinx dx\]\[-\int\limits \sin ^{4}xu ^{2}du=-\int\limits (1-u ^{2})^{2}u ^{2}du=-\int\limits (u ^{2}-2u ^{4}+u ^{6})du\]

OpenStudy (caozeyuan):

I think I did it right

OpenStudy (anonymous):

yeah you did :)

OpenStudy (xapproachesinfinity):

oh something is wrong?

OpenStudy (anonymous):

hm?

OpenStudy (xapproachesinfinity):

i would choose u=sinx du=cosxdx

OpenStudy (xapproachesinfinity):

u^4=sin^4x

OpenStudy (anonymous):

the book went with u=cosx

OpenStudy (caozeyuan):

why?

OpenStudy (xapproachesinfinity):

so int(u^4udu)=int(u^5du)

OpenStudy (anonymous):

but it made everything all negative to fix for it

OpenStudy (xapproachesinfinity):

no no need for that mess realize that cosxdx is du that's wat you need to observe so we need to me u=sinx then du=cosxdx

OpenStudy (xapproachesinfinity):

oh hold on it should be cos^2x not cosx lol

OpenStudy (anonymous):

yeah i think that's why it picked cos as u

OpenStudy (anonymous):

so then it could make u^2 etc

OpenStudy (anonymous):

as opposed to du^2

OpenStudy (xapproachesinfinity):

I'm wrong! i went with cosx not cos^2x his solution is correct

OpenStudy (anonymous):

calc is friggin hard man it hurts my brain

OpenStudy (anonymous):

calc 1 was so easy and then this crap

OpenStudy (xapproachesinfinity):

eh cal2 isn't really that hard, just some attention all there is to it some trig, partial fraction and some other techniques meaning basically dealing with algebra rather than calculus

OpenStudy (anonymous):

yeah its the trig thats killing me

OpenStudy (xapproachesinfinity):

if you got cal1 story you won't have problem with this Calc

OpenStudy (anonymous):

all that fiddling and wiggling crap around trying to make it fit. it's like theres almost no structure to it it always depends on what it is and you gotta fiddle with it

OpenStudy (xapproachesinfinity):

Most students face algebra problem rather than calc namely trig

OpenStudy (anonymous):

<3 algebra. there are so many rules and they're always the same. but with trig it always depends and that kills me

OpenStudy (xapproachesinfinity):

yeah! well you are excepted to have good knowledge on trigs already all they do is form some composition of trig functions and logs exponential and also polynomials

OpenStudy (xapproachesinfinity):

or sometimes product of those six functions!

OpenStudy (xapproachesinfinity):

well you might want to check trig again see the identities to reconnect things because there is no helping it without it! I acually don't remember all trig stuff lol namely secx cscx, cotx! when i studied trig and calc we never paid attention to those but at any rate one can always derive things once you know how to handle cosx sinx

OpenStudy (caozeyuan):

I am doing calc3 which is multi stuff and that is really painful